OpenStudy (anonymous):

it says Solve x^2+6x-16=0... and then it says to type each solution once, use commas to seperate answers needed.

OpenStudy (jamesj):

What are factors of -16 that add up to +6?

OpenStudy (jamesj):

call them a and b. then x^2 + 6x - 16 = (x-a)(x-b) = 0

OpenStudy (anonymous):

heres the numbers without the signs 2 and 8 - where do the signs go?